Evaluate the piecewise defined function at the indicated values. if xs-1 if -1 <S1 x2 + 2x

f(40)
f(-3/2)
f(-1)
f(0)
f(-3)
Answer #1

We want to calculate

f(40), f(-3/2), f(-1), f(0), f(-3).

Now, since 40 is greater than 1. And for x>1, f is defined as:

f(x) = -1

Hence , f(40)= -1.

Since, -3/2 is less than -1 , and for x less than or equal to -1, f is defined as:

f(x) = x​​​​​​​​​​​​​​​​2 + 2x,

Hence f(-3/2) = (-3/2)2 + 2. (-3/2)= -3/4.

Now, for f(-1),

f(-1)= (-1)2 + 2 (-1)= -1,

f(-1)= -1.

Now, since 0 lies between -1 and 1 , and for these values, f is defined as:

f(x)= x ; -1 < x < 1 and for x=1.

Thus, f(0) = 0.

And since -3 < -1, hence

f(-3) = (-3)2 + 2. ( -3) = 9 - 6 = 3,

Hence, f(-3) = 3.